第5章連続確率変数
Chapter Review章の復習
If X has an exponential distribution with mean μ, then the decay parameter is m = , and we write X ∼ Exp(m) where x ≥ 0 and m > 0 . The probability density function of X is f(x) = me-mx (or equivalently . The cumulative distribution function of X is P(X ≤ x) = 1 – e–mx.
Xが平均μの指数分布に従うとき、減衰母数はm = であり、X ∼ Exp(m)と書く。ここでx ≥ 0、m > 0である。Xの確率密度関数はf(x) = me-mx(同じ意味で)となる。Xの累積分布関数はP(X ≤ x) = 1 – e–mxである。
The exponential distribution has the memoryless property, which says that future probabilities do not depend on any past information. Mathematically, it says that P(X > x + k|X > x) = P(X > k).
指数分布には無記憶性がある。これは将来の確率が過去のどの情報にも依存しないことを意味する。数式ではP(X > x + k|X > x) = P(X > k)となる。
If T represents the waiting time between events, and if T ∼ Exp(λ), then the number of events X per unit time follows the Poisson distribution with mean λ. The probability density function of X is . This may be computed using a TI-83, 83+, 84, 84+ calculator with the command poissonpdf(λ, k). The cumulative distribution function P(X ≤ k) may be computed using the TI-83, 83+,84, 84+ calculator with the command poissoncdf(λ, k).
Tが事象の発生間隔を表し、T ∼ Exp(λ)であるなら、単位時間当たりの発生回数Xは平均λのポアソン分布に従う。Xの確率密度関数はである。これはTI-83、83+、84、84+のpoissonpdf(λ, k)コマンドで計算できる。累積分布関数P(X ≤ k)は、TI-83、83+、84、84+のpoissoncdf(λ, k)コマンドで計算できる。
Formula Review公式の復習
Exponential: X ~ Exp(m) where m = the decay parameter
指数分布:X ~ Exp(m)。ここでm = 減衰母数である。
- pdf: f(x) = me(–mx) where x ≥ 0 and m > 0確率密度関数:f(x) = me(–mx)。ここでx ≥ 0、m > 0である。
- cdf: P(X ≤ x) = 1 – e(–mx)累積分布関数:P(X ≤ x) = 1 – e(–mx)
- mean µ =平均μ =
- standard deviation σ = µ標準偏差σ = μ
- percentile k: k =パーセンタイルk:k =
- Additionallyさらに、
- P(X > x) = e(–mx)
- P(a < X < b) = e(–ma) – e(–mb)
- Memoryless Property: P(X > x + k|X > x) = P (X > k)無記憶性:P(X > x + k|X > x) = P(X > k)
- Poisson probability: with mean λポアソン確率:平均λに対し
- k! = k*(k-1)*(k-2)*(k-3)*…3*2*1
References参考文献
Data from the United States Census Bureau.
データ出典:米国国勢調査局。
Data from World Earthquakes, 2013. Available online at http://www.world-earthquakes.com/ (accessed June 11, 2013).
データ出典:World Earthquakes、2013年。http://www.world-earthquakes.com/(2013年6月11日閲覧)。
“No-hitter.” Baseball-Reference.com, 2013. Available online at http://www.baseball-reference.com/bullpen/No-hitter (accessed June 11, 2013).
「ノーヒッター」。Baseball-Reference.com、2013年。http://www.baseball-reference.com/bullpen/No-hitter(2013年6月11日閲覧)。
Zhou, Rick. “Exponential Distribution lecture slides.” Available online at www.public.iastate.edu/~riczw/stat330s11/lecture/lec13.pdf (accessed June 11, 2013).
Zhou, Rick。「指数分布の講義スライド」。www.public.iastate.edu/~riczw/stat330s11/lecture/lec13.pdf(2013年6月11日閲覧)。
Exercises 60–69練習問題 60–69
Use the following information to answer the next ten exercises. A customer service representative must spend different amounts of time with each customer to resolve various concerns. The amount of time spent with each customer can be modeled by the following distribution: X ~ Exp(0.2)
英語のヒント
次の情報を使って、この後の10問に答えよ。顧客サービスの担当者は、さまざまな相談に対処するため、客ごとに異なる時間をかけなければならない。各客にかける時間は、次の分布でモデル化できる。X ~ Exp(0.2)
Exercise 60問題 60
What type of distribution is this?
これはどの種類の分布か。
Exercise 61問題 61
Are outcomes equally likely in this distribution? Why or why not?
この分布では、各結果は同様に確からしいか。理由も述べよ。
Solution解き方
No, outcomes are not equally likely. In this distribution, more people require a little bit of time, and fewer people require a lot of time, so it is more likely that someone will require less time.
同様に確からしくはない。この分布では、少しの時間で済む人が多く、長い時間が必要な人は少ない。そのため、短い時間で済む可能性のほうが高い。
Exercise 62問題 62
What is m? What does it represent?
mはいくつか。何を表すか。
Exercise 63問題 63
What is the mean?
平均はいくつか。
Solution解き方
five
5
Exercise 64問題 64
What is the standard deviation?
標準偏差はいくらか。
Exercise 65問題 65
State the probability density function.
確率密度関数を示せ。
Solution解き方
f(x) = 0.2e-0.2x
Exercise 66問題 66
Graph the distribution.
分布のグラフを描け。
Exercise 67問題 67
Find P(2 < x < 10).
P(2 < x < 10)を求めよ。
Solution解き方
0.5350
Exercise 68問題 68
Find P(x > 6).
P(x > 6)を求めよ。
Exercise 69問題 69
Find the 70th percentile.
第70パーセンタイルを求めよ。
Solution解き方
6.02
Exercises 70–77練習問題 70–77
Use the following information to answer the next seven exercises. A distribution is given as X ~ Exp(0.75).
次の情報を使って、この後の7問に答えよ。分布はX ~ Exp(0.75)である。
Exercise 70問題 70
What is m?
mはいくつか。
Exercise 71問題 71
What is the probability density function?
確率密度関数は何か。
Solution解き方
f(x) = 0.75e-0.75x
Exercise 72問題 72
What is the cumulative distribution function?
累積分布関数は何か。
Exercise 73問題 73
Draw the distribution.
分布を描け。
Exercise 74問題 74
Find P(x < 4).
P(x < 4)を求めよ。
Exercise 75問題 75
Find the 30th percentile.
第30パーセンタイルを求めよ。
Solution解き方
0.4756
Exercise 76問題 76
Find the median.
中央値を求めよ。
Exercise 77問題 77
Which is larger, the mean or the median?
平均と中央値はどちらが大きいか。
Solution解き方
The mean is larger. The mean is , which is greater than 0.9242.
平均のほうが大きい。平均はであり、0.9242より大きい。
Exercises 78–85練習問題 78–85
Use the following information to answer the next 16 exercises. Carbon-14 is a radioactive element with a half-life of about 5,730 years. Carbon-14 is said to decay exponentially. The decay rate is 0.000121. We start with one gram of carbon-14. We are interested in the time (years) it takes to decay carbon-14.
英語のヒント
次の情報を使って、この後の16問に答えよ。炭素14は半減期が約5,730年の放射性元素である。炭素14は指数関数的に崩壊するといわれる。崩壊率は0.000121である。最初に炭素14が1グラムあるとする。炭素14が崩壊するまでの時間(年)に関心がある。
Exercise 78問題 78
What is being measured here?
ここでは何を測定しているか。
Exercise 79問題 79
Are the data discrete or continuous?
データは離散型か、連続型か。
Solution解き方
continuous
連続型。
Exercise 80問題 80
In words, define the random variable X.
確率変数Xを言葉で定義せよ。
Exercise 81問題 81
What is the decay rate (m)?
崩壊率(m)はいくつか。
Solution解き方
m = 0.000121
Exercise 82問題 82
The distribution for X is ______.
Xの分布は______である。
Exercise 83問題 83
Find the amount (percent of one gram) of carbon-14 lasting less than 5,730 years. This means, find P(x < 5,730).
5,730年未満で崩壊する炭素14の量(1グラムに対する百分率)を求めよ。すなわち、P(x < 5,730)を求めよ。
- Sketch the graph, and shade the area of interest.グラフの概形を描き、求める領域に陰影を付けよ。Figure 5.51図 5.51
- Find the probability. P(x < 5,730) = __________確率を求めよ。P(x < 5,730) = __________
Solution解き方
- Answers may vary.解答は一通りではない。
- P(x < 5,730) = 0.5001
Exercise 84問題 84
Find the percentage of carbon-14 lasting longer than 10,000 years.
10,000年を超えて残る炭素14の百分率を求めよ。
- Sketch the graph, and shade the area of interest.グラフの概形を描き、求める領域に陰影を付けよ。Figure 5.52図 5.52
- Find the probability. P(x > 10,000) = ________確率を求めよ。P(x > 10,000) = ________
Exercise 85問題 85
Thirty percent (30%) of carbon-14 will decay within how many years?
何年以内に炭素14の30%が崩壊するか。
- Sketch the graph, and shade the area of interest.グラフの概形を描き、求める領域に陰影を付けよ。Figure 5.53図 5.53
- Find the value k such that P(x < k) = 0.30.P(x < k) = 0.30となるkの値を求めよ。
Solution解き方
- Answers may vary.解答は一通りではない。
- k = 2947.73

